OpenStudy (anonymous):

What is the slope of a line perpendicular to A (-4, 2) and B(4, 2)? A (-4, 2) and B(4, 2) have a slope of 0, so how can I find the slope to the perpendicular line?

OpenStudy (anonymous):

the slope of a line of A (-4, 2) and B(4, 2) is \[2-2 \over 4+4\] = 0 so have a line perpendicular to it has to be opp and negative of what ever the slope is. so the slope perpendicular to A (-4, 2) and B(4,2) would also be 0.

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (anonymous):

no problem ;)

OpenStudy (kropot72):

Wouldn't the slope be \[-\frac{1}{0}=-\infty\]

OpenStudy (anonymous):

There was no slope.